OpenStudy (anonymous):

Using a directrix of y = 2 and a focus of (3, -4), what quadratic function is created?

OpenStudy (tkhunny):

Have you actually heard the word "directrix" in ANY context other than a parabola?

OpenStudy (anonymous):

nooooooooooope

OpenStudy (anonymous):

but these are my options :T f(x) = one twelfth (x - 3)2 - 1 f(x) = -one sixth (x + 3)2 + 1 f(x) = one sixth (x - 3)2 + 1 f(x) = -one twelfth (x - 3)2 - 1

OpenStudy (tkhunny):

Okay, Focus: (3,-4) and Directrix y = 2, where is the Vertex?

OpenStudy (tkhunny):

That's a lovely guess. How did you decide on that?

OpenStudy (anonymous):

v.v ( i'm wrong huh that's like the axis of symmetry ) i thought about "x"

OpenStudy (imstuck):

The info you have so far looks like this on a graph:

OpenStudy (tkhunny):

I said nothing of correct or incorrect. i asked only where the Vertex is. Where is it?

OpenStudy (imstuck):



OpenStudy (anonymous):

i don't know

OpenStudy (imstuck):

The vertex will be located halfway between the y = 2 and y = -4 line and x = 3, like this:

OpenStudy (imstuck):